Find the total resistance if three resistors are in parallel and again another two more resistors are in parallel again

1
Expert's answer
2022-01-18T10:23:30-0500

Assuming all the 5 resistors have the same resistance,

Let "R" = resistance of each resistor.


Effective resistance of the first three in parallel:


"\frac{1}{R_{eff}}= \frac{1}{R}+\frac{1}{R}+\frac{1}{R}=\frac{3}{R}"


"\implies R_{eff}=\frac{R}{3}"


Effective resistance "R_T" of the two resistance in parallel is give by


"R_T= \frac{R*R}{R+R}=\frac{R}{2}"


The set of 3 resistor and the set of two resistors are in series.

Overall resistance "R_o=\frac{R}{2}+\frac{R}{3}=\frac{5}{6}R"